# One-Shot is Enough: Consolidating Multi-Turn Attacks into Efficient Single-Turn Prompts for LLMs

### v1

@@ Abstract
Despite extensive safety enhancements in large language models (LLMs), multi-turn “jailbreak” conversations crafted by skilled human adversaries can still breach even the most sophisticated guardrails. However, these multi-turn attacks demand considerable manual effort, limiting their scalability. In this work, we introduce a novel approach called Multi-turn-to-Single-turn (M2S) that systematically converts multi-turn jailbreak prompts into single-turn attacks. Specifically, we propose three conversion strategies—Hyphenize, Numberize, and Pythonize—each preserving sequential context yet packaging it in a single query. Our experiments on the Multi-turn Human Jailbreak (MHJ) dataset show that M2S often increases or maintains high Attack Success Rates (ASRs) compared to original multi-turn conversations. Notably, using a StrongREJECT-based evaluation of harmfulness, M2S achieves up to 95.9% ASR on Mistral-7B and outperforms original multi-turn prompts by as much as 17.5% in absolute improvement on GPT-4o. Further analysis reveals that certain adversarial tactics, when consolidated into a single prompt, exploit structural formatting cues to evade standard policy checks. These findings underscore that single-turn attacks—despite being simpler and cheaper to conduct—can be just as potent, if not more, than their multi-turn counterparts. Our findings underscore the urgent need to reevaluate and reinforce LLM safety strategies, given how adversarial queries can be compacted into a single prompt while still retaining sufficient complexity to bypass existing safety measures.
@@
Junwoo Ha1 2111Equal contribution., Hyunjun Kim1 3111Equal contribution., Sangyoon Yu1 4, Haon Park1 4, Ashkan Yousefpour1 4, Yuna Park5 6, Suhyun Kim6222Corresponding author: dr.suhyun.kim@gmail.com 1AIM Intelligence,  2University of Seoul,  3Korea Advanced Institute of Science and Technology, 4Seoul National University,  5Yonsei University,  6Korea Institute of Science and Technology
@@ 1 Introduction
The widespread integration of large language models (LLMs) in both industry and academia has not only demonstrated their vast utility but also driven extensive research into developing robust safety mechanisms and ethical deployment practices Carlini et al. (2021); Kandpal et al. (2024); Lukas et al. (2023); Wei et al. (2023); Wen et al. (2023); Zou et al. (2023). In response to potential misuse, most contemporary LLMs are engineered with safety mechanisms designed to refuse tasks that could lead to illegal or unethical outcomes Bai et al. (2022); Ouyang et al. (2022). Despite these precautions, recent studies have revealed that adversaries can exploit vulnerabilities through so-called "jailbreak" attacks—carefully or unintentionally crafted inputs that bypass built-in safeguards and compel the model to generate harmful content Glaese et al. (2022); Korbak et al. (2023).
@@
Motivated by this trade-off, we propose three simple, rule-based Multi-turn-to-Single-turn (M2S) methods as the first systematic approach to transform multi-turn jailbreak conversations into single-turn prompts. Our M2S methods comprise three formatting strategies—Hyphenize, which converts each turn into a bullet-pointed list; Numberize, which uses numerical indices to preserve the sequential order; and Pythonize, which leverages a code-like structure to encapsulate the entire conversation. Despite their simplicity, these methods effectively preserve the high Attack Success Rate (ASR) characteristic of multi-turn human jailbreaks while harnessing the efficiency and scalability of single-turn jailbreaks. To evaluate our approach, we conducted experiments using the Multi-turn Human Jailbreak (MHJ) dataset Li et al. (2024). We evaluated our three M2S methods using the StrongREJECT evaluatorSouly et al. (2024) anchored by three core metrics:
  • Average StrongREJECT Score: Continuous 0-1 harmfulness scale (1.0 = harmful, 0.0 = safe)
  • ASR (%): ASR based on the threshold($\geq$ 0.25 StrongREJECT Score; threshold validated via F1-optimization with human alignment; see Section 4.3)
@@
2 Related Work: Multi-Turn Human Jailbreaks
@@
Evaluating Jailbreaks. When evaluating model responses to jailbreaks attempts, manual or automated evaluation methods can be used. Many prior benchmarks have relied on binary metrics that credited any policy violation or toxic output as a successful jailbreak Wei et al. (2023); Liu et al. (2024); Yu et al. (2024); Xu et al. (2024); Shah et al. (2023); Zhan et al. (2024); Perez et al. (2022); Shaikh et al. (2023); Deng et al. (2024), potentially overestimating effectiveness when the responses were irrelevant or nonsensical. In contrast, the StrongREJECT automated evaluator quantifies harmfulness on a continuous scale by assessing how effectively a response facilitates illicit intent Souly et al. (2024). This approach has demonstrated high agreement with human judgments, thereby providing a more stringent measure of jailbreak success.
@@ 3 Methodology
Figure 1: Overall Pipeline: (1) Take Original Multi-Turn Conversations, (2) Apply M2S Methods, and (3) Evaluate the Resulting Single-Turn Prompts.
@@ 3.3 Pythonize
for i, (question, response) in enumerate(zip(questions, responses), 1):
@@ 4.3 Evaluation Metrics
Threshold Estimation Procedure. Given that the StrongREJECT judge produces continuous harmfulness scores, it is necessary to define a threshold to convert these scores into binary labels for determining whether a response violates safety policies. We first conducted a human evaluation on the final responses from each multi-turn conversation in the Multi-turn Human Jailbreak (MHJ) dataset***https://anonymous.4open.science/r/m2s_dataset-FE80/ where two AI safety experts reached a consensus on each response’s harmfulness. These expert-labeled judgments served as our primary ground truth. By searching for the threshold that maximizes the F1 score, we initially identified an optimal threshold of 0.25.
@@ 5 Results
In this section, we compare the effectiveness of our M2S (Multi-turn-to-Single-turn) conversion methods against the original multi-turn jailbreaks. We focus on three primary dimensions: (i) Attack Success Rate (ASR), Harmfulness, Guardrail Bypass Rate (Tables 12), (ii) Method Adoption Frequencies (Table 3), and (iii) Tactic-Specific Behavior. (Tables 4, 5, and 6). Our findings show that single-turn prompts—carefully constructed from multi-turn jailbreak conversations—can achieve comparable or even higher harmfulness levels and ASRs, despite losing the iterative back-and-forth characteristic of true multi-turn interactions.
@@
Model Method Adoption Frequency (%)
@@ 5.3 Analysis of Tactic-Specific Performance
We turn to the tactic-level analysis, which separates prompts into three outcome categories: Score Increase, Consistent High-Score, and Score Drop. Our findings indicate that certain adversarial tactics—such as Irrelevant Distractor Instructions—gain potency when moved to single-turn format, while others—like Instructing the Model to Continue from the Refusal—appear to rely on multi-turn structure to be fully effective. This has implications for both red-teamers (who can target tactics that flourish in single-turn prompts) and model developers (who should address these newly revealed vulnerabilities). Detailed results in Appendix (Tables 45 and  6).
@@ 6 Conclusion
Our systematic investigation demonstrates that Multi-turn-to-Single-turn (M2S) conversion methods effectively bridge the gap between multi-turn jailbreaks and single-turn jailbreaks. By reformulating iterative adversarial dialogues into structured single-turn prompts—via Hyphenize, Numberize, or Pythonize techniques—we achieve higher attack success rates (ASRs) and enhanced harmfulness scores compared to original multi-turn interactions. The Pythonize method emerges as particularly potent for code-savvy models, while Hyphenize excels in models favoring hierarchical formatting, revealing architecture-dependent parsing vulnerabilities.
@@
7 Limitation and Future Work
Our exploration of Multi-Turn-to-Single-Turn (M2S) conversion methods for jailbreak attacks reveals promising avenues for balancing adversarial potency with operational efficiency. Nevertheless, our work is subject to several limitations that point toward valuable future research directions.
Methodological Constraints in Tactical Adaptation.
In our current evaluation, we derive the M2S performance for each multi-turn conversation by selecting the highest harmfulness score among the three conversion variants—Pythonize, Hyphenize, and Numberize. This best-case metric represents the maximum adversarial performance achievable via M2S; however, it does not account for the specific adversarial tactics employed in each conversation. An automated system that selects the optimal M2S method based on extracted tactic profiles may further improve Attack Success Rates and harmfulness.
Absence of an End-to-End Automated Framework.
Our experimental design currently relies on a semi-automated process, wherein multi-turn jailbreak conversations are manually processed to extract adversarial tactics and then evaluated using our M2S methods. We did not implement an automated framework that integrates tactic extraction, method selection, and subsequent evaluation using the StrongREJECT evaluator. An complete end-to-end system would automatically build the multi-turn dataset, dynamically select the optimal M2S conversion for each conversation based on its tactical profile, and evaluate the resulting responses in a fully automated manner. Such a framework would significantly reduce manual intervention, enhance reproducibility, and promote community-driven improvements in adversarial evaluation.
In summary, while our study demonstrates that M2S conversion methods can effectively bridge the gap between multi-turn effectiveness and single-turn efficiency, future work should focus on developing a tactic-aware selection process and a fully automated, open-source framework. These enhancements are expected to yield more robust performance metrics and provide deeper insights into the interplay between adversarial tactics and conversion efficacy.
Appendix A Appendix

### v2

@@ Abstract
We introduce a novel framework for consolidating multi-turn adversarial “jailbreak” prompts into single-turn queries, significantly reducing the manual overhead required for adversarial testing of large language models (LLMs). While multi-turn human jailbreaks have been shown to yield high attack success rates (ASRs), they demand considerable human effort and time. Our proposed Multi-turn-to-Single-turn (M2S) methods—Hyphenize, Numberize, and Pythonize—systematically reformat multi-turn dialogues into structured single-turn prompts. Despite eliminating iterative back-and-forth interactions, these reformatted prompts preserve and often enhance adversarial potency: in extensive evaluations on the Multi-turn Human Jailbreak (MHJ) dataset, M2S methods yield ASRs ranging from 70.6% to 95.9% across various state-of-the-art LLMs. Remarkably, our single-turn prompts outperform the original multi-turn attacks by up to 17.5% in absolute ASR, while reducing token usage by more than half on average. Further analyses reveal that embedding malicious requests in enumerated or code-like structures exploits “contextual blindness,” undermining both native guardrails and external input-output safeguards. By consolidating multi-turn conversations into efficient single-turn prompts, our M2S framework provides a powerful tool for large-scale red-teaming and exposes critical vulnerabilities in contemporary LLM defenses. All code, data, and conversion prompts are available for reproducibility and further investigations: https://github.com/Junuha/M2S_DATA
@@
Junwoo Ha1 2111Equal contribution., Hyunjun Kim1 3111Equal contribution., Sangyoon Yu1 4, Haon Park1 4, Ashkan Yousefpour1 4, Yuna Park5 6, Suhyun Kim7222Corresponding author: dr.suhyun.kim@gmail.com 1AIM Intelligence,  2University of Seoul,  3Korea Advanced Institute of Science and Technology, 4Seoul National University,  5Yonsei University,  6Korea Institute of Science and Technology,  7Kyung Hee University
@@ 1 Introduction
The widespread integration of large language models (LLMs) in both industry and academia has not only demonstrated their vast utility but also driven extensive research into developing robust safety mechanisms and ethical deployment practices Carlini et al. (2021); Kandpal et al. (2024); Lukas et al. (2023); Wei et al. (2023); Wen et al. (2023); Zou et al. (2023). In response to potential misuse, most contemporary LLMs are engineered with safety mechanisms designed to refuse tasks that could lead to illegal or unethical outcomes Bai et al. (2022); Ouyang et al. (2022). Despite these precautions, recent studies have revealed that adversaries can exploit vulnerabilities through so-called “jailbreak” attacks—carefully or unintentionally crafted inputs that bypass built-in safeguards and compel the model to generate harmful content Glaese et al. (2022); Korbak et al. (2023).
@@
Motivated by this trade-off, we propose three simple, rule-based Multi-turn-to-Single-turn (M2S) methods as the first systematic approach to transform multi-turn jailbreak conversations into single-turn prompts. Our M2S methods comprise three formatting strategies—Hyphenize, which converts each turn into a bullet-pointed list; Numberize, which uses numerical indices to preserve the sequential order; and Pythonize, which leverages a code-like structure to encapsulate the entire conversation. Despite their simplicity, these methods effectively preserve the high Attack Success Rate (ASR) characteristic of multi-turn human jailbreaks while harnessing the efficiency and scalability of single-turn jailbreaks. To evaluate our approach, we conducted experiments using the Multi-turn Human Jailbreak (MHJ) dataset Li et al. (2024). Additionally, the entire dataset and our single-turn conversion prompts (M2S) are publicly available at https://github.com/Junuha/M2S_DATA, enabling researchers to reproduce and extend our findings. We evaluated our three M2S methods using the StrongREJECT evaluatorSouly et al. (2024) anchored by three core metrics:
  • Average StrongREJECT Score: Continuous 0–1 harmfulness scale (1.0 = harmful, 0.0 = safe)
  • ASR (%): ASR based on the threshold ($\geq 0.25$ StrongREJECT Score; threshold validated via F1-optimization with human alignment; see Section 4.3)
@@
2 Related Work
@@
Evaluating Jailbreaks. When evaluating model responses to jailbreaks attempts, manual or automated evaluation methods can be used. Many prior benchmarks have relied on binary metrics that credited any policy violation or toxic output as a successful jailbreak Wei et al. (2023); Liu et al. (2024); Yu et al. (2024b); Xu et al. (2024); Shah et al. (2023); Zhan et al. (2024); Perez et al. (2022); Shaikh et al. (2023); Deng et al. (2024), potentially overestimating effectiveness when the responses were irrelevant or nonsensical. In contrast, the StrongREJECT automated evaluator quantifies harmfulness on a continuous scale by assessing how effectively a response facilitates illicit intent Souly et al. (2024). This approach has demonstrated high agreement with human judgments, thereby providing a more stringent measure of jailbreak success.
@@ 3 Methodology
Figure 1: Overall Pipeline: (1) Take Original Multi-Turn Conversations, (2) Apply M2S Methods.
@@ 3.3 Pythonize
    for i, (question, response) in
    enumerate(zip(questions, responses), 1):
@@ 4.3 Evaluation Metrics
Threshold Estimation Procedure. Given that the StrongREJECT judge produces continuous harmfulness scores, it is necessary to define a threshold to convert these scores into binary labels for determining whether a response violates safety policies. We first conducted a human evaluation on the final responses from each multi-turn conversation in the Multi-turn Human Jailbreak (MHJ) dataset where two AI safety experts reached a consensus on each response’s harmfulness. These expert-labeled judgments served as our primary ground truth. By searching for the threshold that maximizes the F1 score, we initially identified an optimal threshold of 0.25.
@@
4.4 Token Counting Setup
While Attack Success Rates and harmfulness scores constitute our primary evaluation metrics, we additionally measure the token usage to assess the cost and practicality of different jailbreak strategies. Measuring token usage is crucial because it influences both inference cost and the risk of exceeding context windows in practical scenarios. For large-scale adversarial testing, shorter prompts can translate into lower API expenses, but do not necessarily guarantee reduced attack success, as we discuss later in Section 5.2. Following OpenAI’s recommendation, we employ the tiktoken library (o200k_base) to calculate token counts in a manner consistent with GPT-style models. Specifically, we compute the total number of input tokens for each jailbreak prompt in two conditions:
  • Multi-turn Format: We concatenate all user turns and model responses within the conversation up to the final adversarial request. This simulates a typical chat-based interaction and captures the cumulative context length.
  • M2S (Single-turn) Format: We measure the token length of the consolidated single-turn prompt produced by our M2S methods (Pythonize, Hyphenize, Numberize). Since the conversation is flattened, only one prompt is fed to the model.
Note that for M2S, we exclude any intermediate model responses between user turns, as the entire conversation is consolidated into one query. This ensures a consistent comparison with the original multi-turn prompt, where each model response is inherently part of the multi-step dialogue.
In Section 5.2, we present the comparative token statistics for the Multi-turn Human Jailbreak (MHJ) dataset, and Appendix A extends this analysis to two additional multi-turn jailbreak datasets.
@@ 5 Results
In this section, we compare the effectiveness of our M2S (Multi-turn-to-Single-turn) conversion methods against the original multi-turn jailbreaks. We focus on three primary dimensions: (i) Attack Success Rate (ASR), Harmfulness, Guardrail Bypass Rate (Tables 12), (ii) Method Adoption Frequencies (Table 3), and (iii) Tactic-Specific Behavior. (Tables 8, 9, and 10). Our findings show that single-turn prompts—carefully constructed from multi-turn jailbreak conversations—can achieve comparable or even higher harmfulness levels and ASRs, despite losing the iterative back-and-forth characteristic of true multi-turn interactions.
@@
Model Method Adoption Freq (%)
@@ 5.1 Overall Performance
5.2 Token Count Comparison and Analysis
Alongside the attack success metrics presented above, we evaluate how many tokens are consumed by each jailbreak format. Table 4 summarizes the token usage for the Multi-turn Human Jailbreak (MHJ) dataset under both multi-turn and M2S (single-turn) formats. We observe that consolidating multiple conversation turns into a single M2S prompt reduces token requirements from 2732.24 down to 1096.36—a 60% decrease in average length.
Format Avg. Token Count Std. Dev.
Multi-turn 2732.24 728.9
M2S 1096.36 340.2
Table 4: Token usage comparison for the MHJ dataset.
Notably, this substantial drop in token count does not hinder attack efficacy; as shown in Section 5.1, M2S prompts maintain or even improve Attack Success Rates relative to their multi-turn counterparts. In other words, a prompt’s brevity does not necessarily reduce its adversarial potential—shorter prompts can be equally or more potent in bypassing guardrails. From a practical standpoint, fewer tokens mean lower inference cost when using commercial APIs, as well as a reduced risk of exceeding context-length limits. Conversely, defenders must recognize that even more compact prompts can be highly adversarial, underscoring the need for robust safeguards that scrutinize consolidated inputs. Detailed token usage statistics for the other two multi-turn jailbreak datasets are included in Appendix A.
@@ 5.4 Analysis of Tactic-Specific Performance
We turn to the tactic-level analysis, which separates prompts into three outcome categories: Score Increase, Consistent High-Score, and Score Drop. Our findings indicate that certain adversarial tactics—such as Irrelevant Distractor Instructions—gain potency when moved to single-turn format, while others—like Instructing the Model to Continue from the Refusal—appear to rely on multi-turn structure to be fully effective. This has implications for both red-teamers (who can target tactics that flourish in single-turn prompts) and model developers (who should address these newly revealed vulnerabilities). Detailed results in Appendix (Tables 89 and  10).
@@ 5.5 Implications for Red-Teamers and Model Designers
5.6 Data Availability
Alongside the human-annotated labels, we release the complete experimental results for three datasets: our primary MHJ Dataset, SafeMT_ATTACK_600Ren et al. (2024), and the CoSafe-DatasetYu et al. (2024a). For every objective, the archive provides the M2S-converted prompt, the responses of four LLMs (GPT-4o-2024-11-20, Llama-3-70B-chat-hf, Mistral-7B-Instruct-v0.3, and GPT-4o-mini-2024-07-18), their StrongREJECT scores, and token counts (input_tokens, response_tokens, total_tokens).
In total, the repository contains 22,992 prompts (8,592 + 9,600 + 4,800), covering the original multi-turn conversations and three M2S variants, making it immediately useful for reproducibility studies and for benchmarking new defence methods.
All data are available at https://github.com/Junuha/M2S_DATA, and we encourage researchers to make full use of the resource and share follow-up analyses.
@@ 6 Conclusion
Our systematic investigation demonstrates that Multi-turn-to-Single-turn (M2S) conversion methods effectively bridge the gap between multi-turn and single-turn jailbreaks. By reformulating iterative adversarial dialogues into structured single-turn prompts—via Hyphenize, Numberize, or Pythonize techniques—we achieve higher attack success rates (ASRs) and enhanced harmfulness scores compared to original multi-turn interactions. The Pythonize method emerges as particularly potent for code-savvy models, while Hyphenize excels in models favoring hierarchical formatting, revealing architecture-dependent parsing vulnerabilities.
@@
7 Limitations and Future Work
While our Multi-Turn-to-Single-Turn (M2S) approach offers a novel framework for consolidating multi-turn jailbreak prompts into highly effective single-turn attacks, several important limitations remain, pointing to promising directions for future research:
  1. 1.
    Dependence on Select Datasets. Our primary evaluation originally relied on the MHJ dataset, which, though diverse in adversarial tactics, might not capture the full complexity of real-world multi-turn jailbreaks. Although we have extended our experiments to two additional datasets—ATTACK_600Ren et al. (2024) and CoSafeYu et al. (2024a)—these still represent curated benchmarks rather than open-ended, “in-the-wild” attacks. Future studies should expand testing to broader, unstructured scenarios (e.g., user-generated multi-turn dialogues on public platforms), ensuring that M2S generalizes across a more extensive range of adversarial strategies and topic domains.
  2. 2.
    Best-Case Performance Metric. In reporting M2S results, we often highlight the highest Attack Success Rate (ASR) across several formatting variants (e.g., Pythonize, Hyphenize, Numberize). Although this illustrates the upper bound of M2S effectiveness, it does not account for scenarios in which attackers cannot systematically select the optimal formatting strategy. Adversaries often adapt prompts iteratively, and real-time constraints may prevent them from exhaustively testing each M2S variant. Incorporating more realistic “on-the-fly” methods—such as automated tactic classifiers or adaptive prompt-generation agents—would provide a fuller picture of attainable, rather than purely best-case, performance.
  3. 3.
    Limited Automation of the Conversion Pipeline. Currently, we transform multi-turn dialogues into single-turn prompts (Hyphenize/Numberize/Pythonize) using rule-based templates. While these templates are deliberately simple, they require offline processing of the original multi-turn text. A fully automated system that:
    1. (a)
      identifies multi-turn adversarial clues in real time,
    2. (b)
      determines the most suitable single-turn “flattening” strategy,
    3. (c)
      re-injects the aggregated prompt into the LLM without human intervention
    would greatly advance both red-teaming and large-scale safety evaluations. Future work should develop end-to-end pipelines capable of online detection, dynamic formatting, and real-time performance measurement (e.g., with automated judges like StrongREJECT or JailbreakBench).
  4. 4.
    Potential Over-Reliance on a Single Automated Judge. Although StrongREJECT consistently aligns well with human annotations and offers a continuous harmfulness scale, there is still a risk of both false positives (classifying harmless text as harmful) and false negatives (missing subtly harmful content). The fixed threshold ($\mathrm{score}\geq 0.25$) was optimized via F1 scoring on specific labeled data, but different contexts might require alternative calibrations. Future research could triangulate multiple evaluation frameworks—including human raters and alternative automated judges—to confirm that M2S’s adversarial potency is robust across diverse safety metrics.
  5. 5.
    Model-Architecture Variability and Interpretability. Our findings indicate that certain M2S formats (e.g., “Pythonize”) excel with specific model families, whereas others (“Hyphenize”) can dominate on different architectures. This variability suggests that each LLM’s pretraining corpus and system prompts may parse structural cues differently. However, explaining why a given formatting style outperforms another on a given LLM remains difficult due to proprietary or black-box model internals. Deeper interpretability studies—potentially via sequence attributions or hidden-state analyses—could help clarify how enumerated versus code-like structures influence an LLM’s vulnerability to jailbreaks.
  6. 6.
    Real-World Complexity Beyond Token Counts. While our experiments show that M2S can drastically reduce token usage (and thus overhead) compared to multi-turn adversarial interactions, real-world deployments often involve additional system prompts, plugin calls, or reasoning traces that inflate total token counts. In a live chat system, for example, the token savings from M2S might be overshadowed by user-specific context or session-level state tokens. Future investigations should benchmark M2S in full, production-grade environments to gauge the practical cost savings and confirm that condensed single-turn prompts still bypass robust, real-time guardrails.
  7. 7.
    No Integrated Defense Proposals. Our study focuses primarily on revealing a new class of powerful single-turn threats rather than proposing comprehensive defenses. Although we highlight certain weaknesses—such as “contextual blindness” to enumerated or code-structured prompts—closing these gaps will likely require multi-layered approaches (e.g., improved token-level content filtering, heuristic detection of “flattened” prompts, or dynamic context-checking across simulated conversation states). Future lines of work should therefore develop and test defense strategies that operate effectively against M2S-style prompts, combining both static analysis (e.g., scanning for enumerations or code blocks) and runtime monitoring (e.g., verifying internal policy compliance per sub-request).
In summary, while our M2S framework underscores a critical vulnerability in current LLM safeguard designs and demonstrates how multi-turn adversarial dialogues can be “flattened” into potent single-turn prompts, a number of open questions remain. Addressing these—through broader datasets, more realistic best-case modeling, fully automated pipelines, multi-judge evaluation, interpretability research, real-world integration, and focused defensive strategies—will be essential for advancing the field of LLM safety and ensuring robust protection against ever-evolving adversarial prompt engineering.
Ethical Considerations
Our work aims to illuminate critical weaknesses in contemporary large-language-model (LLM) safety mechanisms by introducing methods that consolidate multi-turn adversarial tactics into single-turn prompts. Below we outline the key ethical considerations involved in designing and disseminating this research.
Intended use and potential misuse
The techniques we propose—Hyphenize, Numberize, and Pythonize—can reveal critical vulnerabilities in LLMs, but they could also be misapplied to generate harmful or disallowed content at scale. We release our methods and data to empower researchers and practitioners to identify and remediate these vulnerabilities, not to promote malicious exploitation. Any such use would contravene the spirit and intent of our work.
Data and content warnings
The datasets employed in this study (e.g. Multi-turn Human Jailbreak, Attack_600, CoSafe) inevitably contain adversarial prompts that request harmful or disallowed outputs from LLMs. We explicitly annotate these datasets as containing potentially sensitive, unsafe, or harmful text and strongly discourage their use outside carefully controlled safety evaluations. The data are released solely to facilitate the development and testing of robust defensive strategies.
Responsible release
We have taken care to limit the direct reproduction of explicitly harmful content in our examples and tables, focusing instead on high-level analyses of adversarial prompts and aggregate experiment results. Where possible, we redact or paraphrase content to minimise the risk of exposing sensitive or disallowed information. The consolidated single-turn prompts are made available only for reproducibility and to enable safety researchers to analyse emerging threats.
Broader impact on LLM safety
By demonstrating that single-turn, “flattened” adversarial prompts can sometimes surpass even carefully orchestrated multi-turn attacks, we aim to motivate the development of stronger guard-rail systems. In particular, our findings underscore the need to scrutinise entire blocks of input comprehensively rather than relying on turn-by-turn context checks alone. We hope that shedding light on these vulnerabilities will ultimately lead to more robust alignment and safer deployment of LLMs.
Overall, we adhere to the principle of coordinated disclosure by publishing our findings with clear disclaimers and minimal direct reproduction of harmful text. We hope this work will guide community-wide initiatives aimed at reinforcing LLM safety and mitigating the risks associated with adversarial attacks.
Appendix A Expanded Token Count Analysis
We have conducted a comprehensive assessment using OpenAI’s officially recommended tiktoken library (o200k_base) to better quantify token usage across different multi-turn jailbreak datasets. To ensure broader applicability beyond the MHJ dataset, we also ran additional experiments on two other multi-turn jailbreak datasets: ATTACK_600 and CoSafe. ATTACK_600 comprises especially lengthy multi-turn instructions crafted by advanced red-teamers, while CoSafe focuses on succinct scenario-driven adversarial queries. By including these diverse datasets, we capture a broader spectrum of conversation styles that challenge LLM guardrails from multiple angles.
Dataset Multi-turn Single-turn (M2S)
ATTACK_600 5331.32 830.61
MHJ 2732.24 1096.36
CoSafe 1689.28 469.10
Table 5: Average token counts measured with tiktoken (o200k_base) across three multi-turn jailbreak datasets.
These findings reinforce our assertion that the multi-turn format can accumulate significantly more tokens due to iterative message-passing between user and model. By contrast, a single-turn M2S prompt tends to be considerably shorter—yet can still retain (or even improve) the overall attack effectiveness, as demonstrated in Section 5.
Moreover, according to OpenAI’s official Conversation State guidelines, the total tokens consumed in an actual chat scenario can exceed these figures. This is because system instructions, internal reasoning tokens, or other role-based content may be appended automatically to maintain context across conversation turns. We hope these extended analyses offer a clearer perspective on how multi-turn interactions can inflate token usage, and why even shorter, single-turn prompts can remain highly adversarial.
Model Turn Method ASR (%) Average Score
GPT-4o-2024-11-20 Multi Original 57.3 0.53
Single Hyphenize (M2S) 49.3 (--8.0) 0.38 (--0.15)
Single Numberize (M2S) 49.3 (--8.0) 0.40 (--0.13)
Single Pythonize (M2S) 60.7 (+3.4) 0.56 (+0.03)
Single Ensemble (M2S) 67.0 (+9.7) 0.59 (+0.06)
Llama-3-70B-chat-hf Multi Original 46.0 0.37
Single Hyphenize (M2S) 38.7 (--7.3) 0.27 (--0.10)
Single Numberize (M2S) 39.3 (--6.7) 0.27 (--0.10)
Single Pythonize (M2S) 54.0 (+8.0) 0.42 (+0.05)
Single Ensemble (M2S) 57.0 (+11.0) 0.46 (+0.09)
Mistral-7B-Instruct-v0.3 Multi Original 63.7 0.46
Single Hyphenize (M2S) 75.7 (+12.0) 0.55 (+0.09)
Single Numberize (M2S) 74.3 (+10.6) 0.54 (+0.08)
Single Pythonize (M2S) 76.0 (+12.3) 0.56 (+0.10)
Single Ensemble (M2S) 86.0 (+22.3) 0.69 (+0.23)
GPT-4o-mini-2024-07-18 Multi Original 61.0 0.50
Single Hyphenize (M2S) 64.7 (+3.7) 0.48 (--0.02)
Single Numberize (M2S) 61.0 (+0.0) 0.45 (--0.05)
Single Pythonize (M2S) 72.0 (+11.0) 0.58 (+0.08)
Single Ensemble (M2S) 77.3 (+16.3) 0.64 (+0.14)
Table 6: M2S Performance on CoSafe. Attack-Success Rate (ASR) and Average StrongREJECT Score across models. Parentheses show the delta relative to the corresponding multi-turn baseline.
Appendix B M2S Performance on CoSafe
Below we report the M2S performance on the CoSafe dataset (see Table 6). The Hyphenize, Numberize, and Pythonize methods are applied to convert multi-turn adversarial queries into single-turn prompts. Overall, we observe that certain models see improvements in Attack Success Rate (ASR) with the Pythonize or ensemble approach, while others exhibit marginal decreases when using Hyphenize or Numberize alone. Notably, the “ensemble” metric (i.e., selecting the best result among the three M2S variants for each query) consistently outperforms the original multi-turn baseline.
Appendix C M2S Performance on ATTACK_600
Similarly, we show performance on the ATTACK_600 dataset (Table 7). Despite being relatively long and complex instructions, we again find that our M2S single-turn methods can retain or even boost overall ASR. The “Pythonize” variant often yields a substantial harmfulness increase, while “Hyphenize” and “Numberize” can still surpass the multi-turn baseline in some cases. Once more, the ensemble scenario demonstrates that consolidating the most effective single-turn result can provide a large improvement over the original multi-turn prompts.
Model Turn Method ASR (%) Average Score
GPT-4o-2024-11-20 Multi Original 71.5 0.61
Single Hyphenize (M2S) 69.0 (--2.5) 0.54 (--0.07)
Single Numberize (M2S) 67.8 (--3.7) 0.54 (--0.07)
Single Pythonize (M2S) 76.8 (+5.3) 0.69 (+0.08)
Single Ensemble (M2S) 91.0 (+19.5) 0.81 (+0.20)
Llama-3-70B-chat-hf Multi Original 63.8 0.42
Single Hyphenize (M2S) 71.0 (+7.2) 0.50 (+0.08)
Single Numberize (M2S) 63.8 (+0.0) 0.46 (+0.04)
Single Pythonize (M2S) 71.7 (+7.9) 0.46 (+0.04)
Single Ensemble (M2S) 92.2 (+28.4) 0.69 (+0.27)
Mistral-7B-Instruct-v0.3 Multi Original 73.2 0.42
Single Hyphenize (M2S) 80.2 (+7.0) 0.55 (+0.13)
Single Numberize (M2S) 80.5 (+7.3) 0.55 (+0.13)
Single Pythonize (M2S) 82.3 (+9.1) 0.54 (+0.12)
Single Ensemble (M2S) 96.5 (+23.3) 0.71 (+0.29)
GPT-4o-mini-2024-07-18 Multi Original 65.0 0.48
Single Hyphenize (M2S) 78.2 (+13.2) 0.56 (+0.08)
Single Numberize (M2S) 75.3 (+10.3) 0.53 (+0.05)
Single Pythonize (M2S) 74.2 (+9.2) 0.48 (+0.00)
Single Ensemble (M2S) 95.7 (+30.7) 0.72 (+0.24)
Table 7: M2S Performance on ATTACK_600. ASR and Average StrongREJECT Score for each model. Parentheses denote the change relative to the multi-turn baseline.
In summary, these additional datasets illustrate the generality of the M2S approach. Although multi-turn jailbreaks typically accumulate lengthy token contexts and can be extremely effective, flattening them into single-turn formats (while preserving the structural manipulations) can maintain or even enhance adversarial potency, often at a fraction of the token usage.
Appendix D Extended Tactic-Level Analysis
In addition to the primary evaluation metrics, we analyze adversarial tactics and how their StrongREJECT scores vary when transitioning from multi-turn to single-turn prompts. Specifically, we identify three types of shifts in harmfulness:
  • Score Increase Tactics: Jailbreak methods that become more harmful (from near 0 to $\geq 0.75$) when reformatted into a single-turn prompt.
  • Consistent High-Score Tactics: Methods that maintain a consistently high StrongREJECT score ($\geq 0.75$) across both multi-turn and single-turn scenarios.
  • Score Drop Tactics: Jailbreak methods that lose effectiveness (from $\geq 0.75$ to 0) when converted to single-turn formatting.